Assume both portfolios A and B are well diversified, that E(rA) = 12.8% and E(rB) = 14.0%. If the economy has only one factor, and βA = 1 while βB = 1.2, what must be the risk-free rate?
Answer #1

Expected return=risk free rate+beta*(market rate-risk free rate)

12.8=Rf+1*(Rm-Rf)

Rm=12.8%

Also:

14=Rf+1.2*(Rm-Rf)

14=1.2*Rm-0.2Rf

14=(1.2*12.8)-0.2Rf

Rf=(15.36-14)/0.2

=6.8%=risk free rate